Hollywood Shoes would like to maintain their cash account at a minimum level of $64,000, but expect the standard deviation in net daily cash flows to be $5,400; the effective annual rate on marketable securities to be 7.00 percent per year; and the trading cost per sale or purchase of marketable securities to be $240 per transaction. What will be their optimal cash return point? (Round your answer to 2 decimal places.)
Answer #1

iday = (1 + r)^(n/365) - 1 = 1.07^(1/365) - 1 = 1.000185383 - 1 = 0.000185383

And the optimal cash return point will be equal to:

Z* = [{3 x Cost per transaction x Variance of Daily Cash Flow} / (4 x Rate of Return on Marketable Securities per day}](1/3) + Lower Limit Control

= [3($240)($5,400)2} / {(4 x 0.000185383)}](1/3) + $64,000

= $30,478.72 + $64,000 = $94,478.72